OpenStudy (anonymous):

If two lines are parallel and the equation of one line is 2x-3y=0, does the other line increase or decrease from left to right? A. Increases B. Decreases C. The line is horizontal D. The line is vertical E. None of the above.

OpenStudy (hoblos):

two parallel lines have same slope.. so we must first get the slope of the line 2x-3y=0 we ca do that by transforming the equation into slope intercept form which has a general form of: y=ax+b where a is the slope so here 2x-3y=0 2y=3x y=(3/2)x thus the slope is 3/2 which is positive a line with positive slope increases from left to right thus the answer is A

OpenStudy (anonymous):

thanks love <3

OpenStudy (hoblos):

glad to help :))